# A Reactive Grasping Framework for Multi-DoF Grippers via Task Space Velocity Fields and Joint Space QP

### v1

@@ Abstract
We present a fast and reactive grasping framework for multi-DoF grippers that combines task-space velocity fields with joint-space Quadratic Program (QP) in a hierarchical structure. Reactive, collision-free global motion planning is particularly challenging for high-DoF systems, as simultaneous increases in state dimensionality and planning horizon trigger a combinatorial explosion of the search space, making real-time planning intractable. To address this, we plan globally in a lower-dimensional task space – such as fingertip positions – and track locally in the full joint space while enforcing all constraints. This approach is realized by constructing velocity fields in multiple task-space coordinates (or, in some cases, a subset of joint coordinates) and solving a weighted joint-space QP to compute joint velocities that track these fields with appropriately assigned priorities. Through simulation experiments with privileged knowledge and real-world tests using the recent pose-tracking algorithm FoundationPose [1], we verify that our method enables high-DoF arm–hand systems to perform real-time, collision-free reaching motions while adapting to dynamic environments and external disturbances. Project page: https://reactivegrasp.github.io.
@@ I Introduction
As robotic systems move toward unstructured and dynamic environments, the ability to adapt in real-time – referred to as reactive control – has become a key capability. Among core manipulation skills, grasping is one of the most fundamental and widely studied. Nonetheless, reactive grasping remains a challenging problem, particularly for multi-DoF grippers, which have received relatively less attention compared to parallel-jaw grippers [2, 3, 4]. A few exceptions explore reinforcement learning approaches [5, 6]; however, these are limited to single-object settings without obstacles and primarily top-down reaching motions. Their applicability to cluttered scenes or more diverse grasp-reaching directions remains unclear.
@@
For reactive grasping with multi-DoF grippers, these methods can be applied in either task space (e.g., at the level of individual links) or joint space, but both present challenges. In task space, multiple potential fields or dynamical systems are usually combined by weighting control forces projected into joint space via the Jacobian. This often results in (i) tedious weight tuning, (ii) poor representation of highly concave obstacles (e.g., wine glasses) when using structured shape abstractions such as tree-of-stars, and (iii) combined fields that lack convergence guarantees and are prone to local minima. In joint space, collision constraints are typically highly non-convex and unstructured, making them difficult to approximate with compact geometric abstractions and limiting the applicability of reactive modulation approaches.
@@
In this paper, we present a highly reactive grasping framework for multi-DoF grippers that avoids the coupled growth of state and time-horizon dimensionality. The key idea is to plan globally in a lower-dimensional task space – keeping the state dimension fixed and low – with a subset of constraints, and to track locally in full joint space – keeping the time horizon fixed and short – while considering all constraints. This is implemented as a two-layer hierarchical control framework.
@@ III-B Real Robot Experiments
An episode is deemed successful if the robot lifts the object. Collisions with the target object are not considered failures, while collisions with other obstacles are (except incidental contacts with unmodeled elements such as cables). Trials were excluded if failures were clearly due to hardware issues – e.g., communication errors, EMI, or motor failures – or if FoundationPose [1] completely lost the object pose, primarily from occlusions. In some cases, the robot could in principle reach the targets but failed to move further due to backlash or stiction in the wrist or pulley–cable mechanisms of our custom arm and hand – for example, stopping even when the fingertips were close to the targets with a collision-free path remaining and no joint limits or constraints active. Such cases, which would not occur with conventional arms, were counted as successes to isolate algorithmic performance from mechanical limitations.
Figure 10: An example of random disturbances, such as external pushes to the robot or shifts in the object’s position.
First, we perform a single-object grasping task without additional obstacles, under external disturbances. Disturbances are applied either to the robot (Fig. 10) or to the object (Fig. 1) after the robot begins moving, at least twice per trial. Table II reports success rates over 5 trials for 4 objects. Grasps were successful in most trials. Although the robot fingers occasionally collided with the target object, the robot generally continued to track the object and complete the grasp. A failure was observed with the tape object, where the robot prematurely closed the gripper while the tape was still rolling, resulting in a missed grasp.
TABLE II: Grasp success rates under random disturbances (successes out of 5 trials).
Bowl Cup Tape Wine Glass
5/5 5/5 4/5 5/5
TABLE III: Grasp success rates for the bowl with varying numbers of obstacles (successes out of 10 trials).
@@
Second, we evaluate grasping performance in the presence of obstacles, as illustrated in Fig. 11. The target object and obstacles are randomly placed, and the reaching controller is executed without external disturbances. Table III reports the success rates over 10 trials for the cases with one, two, and three obstacles. Fig. 11 illustrates a successful case. The executed fingertip paths are shown with inter-fingertip segments (Left), alongside examples of planned fingertip paths (Right). From the right figure, two observations can be made: (i) the target grasp positions change as the motion progresses, and (ii) the planned and executed paths do not coincide exactly. The first observation is expected, since grasp targets are updated online based on the current fingertip positions. The second arises because the low-level joint-space QP must enforce constraints and track additional velocity objectives at the same time, making exact tracking of the planned fingertip velocities inherently impossible. As shown in the figure, the planned fingertip paths nevertheless serve to continuously guide the fingertip motion toward the targets in real time, while avoiding collisions with the obstacles and the object.
@@ IV Discussion and Conclusion
We have proposed a hierarchical control framework for reactive grasping that is robust to external disturbances, where the higher layer defines task-space (or joint-subspace) velocity targets and the lower layer tracks them under full joint-level constraints with appropriate weighting. By leveraging global path planning to generate high-level fingertip velocity targets, the framework enables robust grasping of highly concave objects, such as wine glasses and bowls, even in the presence of obstacles, thus advancing reactive manipulation in cluttered and dynamic environments.
Our framework can be further enhanced in two directions. First, it assumes a set of pre-defined grasp points, which limits adaptability to novel objects; this could be addressed by incorporating learning-based methods to predict favorable grasp points online. Second, in cluttered environments with multiple obstacles, the system may become trapped in local minima when only fingertip velocity targets are considered (see Fig. 12); incorporating additional velocity targets at the elbow or other joints, potentially through lower-frequency joint-level optimization as in [41], represents a promising extension.

### v2

# Hierarchical Reactive Grasping via Task-Space Velocity Fields and Joint-Space Quadratic Programming

@@ Abstract
We present a fast and reactive grasping framework that combines task-space velocity fields with joint-space Quadratic Program (QP) in a hierarchical structure. Reactive, collision-free global motion planning is particularly challenging for high-DoF systems, as simultaneous increases in state dimensionality and planning horizon trigger a combinatorial explosion of the search space, making real-time planning intractable. To address this, we plan globally in a lower-dimensional task space – such as fingertip positions – and track locally in the full joint space while enforcing all constraints. This approach is realized by constructing velocity fields in multiple task-space coordinates (or, in some cases, a subset of joint coordinates) and solving a weighted joint-space QP to compute joint velocities that track these fields with appropriately assigned priorities. Through simulation experiments and real-world tests using the recent pose-tracking algorithm FoundationPose [1], we verify that our method enables high-DoF arm–hand systems to perform real-time, collision-free reaching motions while adapting to dynamic environments and external disturbances. Project page: https://reactivegrasp.github.io.
@@ I Introduction
As robotic systems move toward unstructured and dynamic environments, the ability to adapt in real-time – referred to as reactive control – has become a key capability. Among core manipulation skills, grasping is one of the most fundamental and widely studied. Nonetheless, reactive grasping remains a challenging problem, particularly for high-dimensional systems, e.g., with multi-DoF grippers.
While recent work [2, 3] has advanced real-time grasp-pose generation and tracking, the complementary challenge of collision-free reactive motion planning has received less attention. RL-based approaches [4, 5, 6] are often restricted to single-object, obstacle-free, top-down reaches, which limits their generalizability to out-of-distribution scenarios. In contrast, we propose a model-based real-time planning and control framework that generalizes naturally across clutter and diverse grasping directions.
@@
For reactive grasping, these methods can be applied in either task space (e.g., at the level of individual links) or joint space, but both present challenges. In task space, multiple potential fields or dynamical systems are usually combined by weighting control forces projected into joint space via the Jacobian. This often results in (i) tedious weight tuning, (ii) poor representation of highly concave obstacles (e.g., wine glasses) when using structured shape abstractions such as tree-of-stars, and (iii) combined fields that lack convergence guarantees and are prone to local minima. In joint space, collision constraints are typically highly non-convex and unstructured, making them difficult to approximate with compact geometric abstractions and limiting the applicability of reactive modulation approaches.
@@
In this paper, we present a highly reactive grasping framework for high-dimensional arm-hand systems that avoids the coupled growth of state and time-horizon dimensionality. The key idea is to plan globally in a lower-dimensional task space – keeping the state dimension fixed and low – with a subset of constraints, and to track locally in full joint space – keeping the time horizon fixed and short – while considering all constraints. This is implemented as a two-layer hierarchical control framework.
@@ III-B Real Robot Experiments
An episode is deemed successful if the robot lifts the object. Collisions with the target object are not considered failures, while collisions with other obstacles are (except incidental contacts with unmodeled elements such as cables). Trials were excluded if failures were clearly due to hardware issues – e.g., motor failures – or if FoundationPose [1] completely lost the object pose, primarily from occlusions, to isolate algorithmic performance from mechanical limitations and perception failures.
Figure 10: Examples of strong disturbances (large deviations in object position and strong pushes to the robot).
First, we evaluate single-object grasping without additional obstacles, under strong disturbances applied either to the robot or to the object (see Fig. 10). These disturbances are introduced after the robot has begun moving and occur multiple times per trial. Specifically, while conducting strictly controlled experiments is challenging, we constrain the disturbances to remain within reasonable bounds: (i) the objects must stay within the robot’s workspace, (ii) they must not be moved into ungraspable configurations (e.g., an upside-down bowl), and (iii) their motion must remain within the speed range that FoundationPose can reliably track (see supplementary videos for examples). We performed five trials per object. The bowl, cup, and wine glass tasks were all successful, whereas the tape task resulted in one failure. Although the robot fingers occasionally collided with the target object due to tracking errors, the system was generally able to recover and complete the grasp. The failure with the tape object – where the gripper closed prematurely while the tape was still rolling, resulting in a missed grasp – was an expected outcome, as no prediction module for the object’s future position was used.
TABLE II: Grasp success rates for the bowl with varying numbers of obstacles (successes out of 10 trials).
@@
Second, we evaluate grasping performance in the presence of obstacles, as illustrated in Fig. 11. The target object and obstacles are randomly placed, and the reaching controller is executed without external disturbances. Table II reports the success rates over 10 trials for the cases with one, two, and three obstacles. Fig. 11 illustrates a successful case. The executed fingertip paths are shown with inter-fingertip segments (Left), alongside examples of planned fingertip paths (Right). From the right figure, two observations can be made: (i) the target grasp positions change as the motion progresses, and (ii) the planned and executed paths do not coincide exactly. The first observation is expected, since grasp targets are updated online based on the current fingertip positions. The second arises because the low-level joint-space QP must enforce constraints and track additional velocity objectives at the same time, making exact tracking of the planned fingertip velocities inherently impossible. As shown in the figure, the planned fingertip paths nevertheless serve to continuously guide the fingertip motion toward the targets in real time, while avoiding collisions with the obstacles and the object.
@@ IV Discussion and Conclusion
We have proposed a hierarchical control framework for reactive grasping that is robust to external disturbances, where the higher layer defines task-space (or joint-subspace) velocity targets and the lower layer tracks them under full joint-level constraints with appropriate weighting. The framework enables robust grasping of highly concave objects, such as wine glasses and bowls, even in the presence of obstacles, thus advancing reactive manipulation in cluttered and dynamic environments.
Our framework can be improved in two directions. First, it relies on pre-defined grasp points, which limits adaptability to novel objects; this could be addressed by incorporating learning-based methods to predict grasp points online. Second, in cluttered environments with multiple obstacles, the system may become trapped in local minima when only fingertip velocity targets are considered (see Fig. 12); adding velocity targets at the elbow or other joints, potentially through lower-frequency joint-level optimization as in [41], is a promising extension.